How many liters of water must be added to 5 liters of a 30% acid solution to obtain a 20% acid solution?
Found 2 solutions by Paul, checkley71:
Answer by Paul(988)

EQUATION:
5(0.3)=0.2(x+5)
1.5=1+0.2x
0.2x=.5
x=2.5
Hence, about 2.5L should be added.
Paul.

Answer by checkley71(8403)

5*.3=.2(5+x) or 1.5=1+.2x or .5=.2x or x=5/.2 or x=2.5 liters of water are added to 5 liters which gives us 7.5 liters of solution witl 1.5 liters of acid which calculates to 1.5/7.5=.20 or 20%